OpenStudy (calculusxy):

Can you please help me solve for a: a - (1 - 1/3)=4 1/3 @Mertsj

OpenStudy (calculusxy):

@Mertsj

OpenStudy (anonymous):

first use the association law

OpenStudy (calculusxy):

what is that law?

OpenStudy (calculusxy):

@satellite73

OpenStudy (mertsj):

I don't know how you solve for a because that is not an equation. You could simplify the expression

OpenStudy (calculusxy):

Yeah simplify

OpenStudy (mertsj):

1-1/3=2/3 a-(2/3)=a-2/3

OpenStudy (calculusxy):

\[u \emptyset v = u - (1- \frac{ 1 }{ v }\] \[a \emptyset 3 = 4 \frac{ 1 }{ 3 }\] what does a equal to?

OpenStudy (mertsj):

I do not understand that notation. I have never seen it before and don't know what it means. Can you find it online somewhere so I can read about it?

OpenStudy (mertsj):

Yes I know the notation for the empty set but I have never seen a variable next to that symbol which typically would mean multiply but it makes no sense. So I am thinking you didn't post the problem completely.